# Community Report from the 2025 SNOLAB Future Projects Workshop

### v1

@@ 3.2 The China Jinping Underground Laboratory
Two planned-to-be-refilled auxiliary tunnels after the hydropower station were selected for the construction of CJPL-II. The total volume of CJPL-II is about 300 km3 with four main halls of 14 m (width) 14 m (height) $\times$ 130 m (length) shown in Figure 4. A large pure water vessel is built in hall B2 with a size of 27 m $\times$ 15 m $\times$ 13 m and is used to shield from environmental radioactivity. A large liquid nitrogen tank, located in hall C1 with a diameter of 13 m and height of 13 m, is used to provide shielding and a cryogenic environment for germanium-based experiments. A low background counting facility located in hall C2 consists of an ultra-low background germanium spectrometer with detection limit of 10 $\mu$Bq/kg chen_background_2025 and fifteen low background germanium spectrometers with detection limit of 0.1 mBq/kg ma_status_2021 .
@@ 4.2 HeLIOS – Detecting Gravitational Waves and Dark Matter with Superfluid Helium
Co-Authors: M. Hirschel and J. P. Davis
@@ 4.3 Liquid-Noble Bubble Chambers
To fully understand the effect of this reduction in electron recoil nucleation probability, the Scintillating Bubble Chamber (SBC) collaboration is building two functionally-identical detectors adopting a buffer-free chamber design bressler2019buffer . The targeted operation (130 K and 30 psi) would allow for a 100 eV nuclear recoil threshold sufficient to probe DM masses in the GeV/$c^{2}$ mass range Alfonso-Pita:2023frp ; Alfonso-Pita:2022akn . In these detectors, 10 kg of liquid argon, doped with xenon to act as a wavelength shifter, is contained in a set of nested, synthetic quartz vessels, and further housed in a pressure vessel, shown in the left of Fig. 6. Thermomechanical regulation is provided by liquid nitrogen thermosyphons for cooling and an external hydraulic piston for pressure control. The first physics-scale detector, SBC-LAr10, will be commissioned this summer in the MINOS tunnel at Fermilab for engineering and calibration studies and in parallel, assembly at SNOLAB for the low-background DM search detector will begin.
@@
Looking forward to the future, a 1-tonne-yr exposure would allow for a DM search down to the boundary of the argon neutrino fog from 8B solar neutrinos, shown in the right of Fig. 6. A 1 tonne superheated liquid argon volume at 130 K would need dimensions of approximately 1-m-diameter by 1.4-m-high. This volume needs to be pressurized to $\sim 350$ psi after a bubble nucleation, and therefore a pressure vessel capable of housing this fluid would require $\lesssim 0.01$ mBq/kg of primordial isotopes to generate $<1$ background event/yr from fast neutrons. While these are small activities, they may not be so far off from some upper limits, emphasizing the need for precision assay methods in the future.
Critically, synthetic quartz vessels, which are traditionally used to contain the superheated liquid, cannot be produced with active volumes greater than 250 L. This provides an opportunity to investigate other detector materials, such as electropolished metals and acrylic, both of which have been demonstrated in small-scale test chambers Broerman:2022jtj . To limit deadtime from repressurizing the detector after a bubble event ($\sim$30 s/event), surface rates from radiogenic and roughness features would need to be kept $<$5 nBq/cm2 to result in $<1$ event/hour. For external shielding, a 9-m-diameter water tank would be sufficient to reduce the fast neutron and Thomson scattering backgrounds from the cavern walls to $<1$ event/year each.
@@ 4.4 Cryogenic Solid-State Detectors: SuperCDMS and Beyond
is a plan to scale up the HVeV concept by combining many detectors, weighing a few grams each, into honeycomb-shaped layers with $\sim 10$ cm faces (see Fig. 7). Deep Reactive Ion Etching would be employed to create trenches in each wafer of crystal material to form the honeycomb. This would leave a small “bridge” for thermal conduction between hexagonal areas in the honeycomb; the smooth dicing would reduce sidewall charge injection. Stress-free mounting would be employed, with the goal of mitigating the low energy excess by reducing events caused by crystal strain. Multiplexing readout with $\mu$MUX, with multi-stage cryogenic amplifiers, would be required. Multiple wafers would be stacked on top of each other, for a total payload of $\mathcal{O}(10)~{}\mathrm{kg}$, with the wafers serving as active vetoes for each other. The energy threshold goal is at the eV level. Recent RF readout upgrades at CUTE make this facility an ideal host.
@@
    There is the possibility of sharing infrastructure with other future cryogenic solid-state detector projects, e.g. Globolo (see Section 5), and synergy with TESSERACT (see Section 3.3.3).
@@ 4.6 Novel Small-Scale Experiments for Light Dark Matter
Co-Authors: Y. Kahn, B. Roach, N. Hoch, A. Williams, E. Marrufo Villalpando, D. Baxter, A. Drlica-Wagner, P. Abbamonte, D. Balut, J. Oh, L. Thompson, D. Johnson, D. Freedman, L. Winslow, C. Blanco, A. Radick, B. Lillard, C. Gao, J. Schütte-Engel, W. Halperin, M. Nguyen, J.W. Scott, and J. Foster
@@
5 Bolometric Detectors: Precision Beta Decay and Future Neutrinoless Double Beta Decay Searches
@@ 6.2 SNO+ Tellurium Double Beta Decay
SNO+ will be moving forward with the $0\nu\beta\beta$ phase of the experiment. Tellurium will be added to the detector, starting in 2026, in the form of tellurium-butanediol complexes dissolved in the SNO+ liquid scintillator. The SNO+ $0\nu\beta\beta$ search starts after the current pure scintillator phase has completed its main physics program, and after test batches have fully commissioned the two underground tellurium process plants (purification and synthesis). In the first instance, 0.5% Te, by weight, will be the target loading. This amounts to 1.33 tonnes of the isotope 130Te in the detector, a large quantity due to its high natural abundance of 34%. A proposal has been submitted to the CFI 2025 Innovation Fund competition. If it is successful, it would enable SNO+ to triple the amount of Te loaded in the detector to 1.5% Te, reaching 4 tonnes of 130Te deployed for the neutrinoless double beta decay search. Figure 8 illustrates the sensitivity that SNO+ 0.5% and SNO+ 1.5% can reach in comparison to other existing, near-future and planned neutrinoless double beta decay experiments.
Figure 8: Comparison of neutrinoless double beta decay sensitivities for current (Section 3.3), near-term and future proposed (Sections 3.3 and 6.6) experiments. Sensitivity is expressed as $m_{\beta\beta}$. Different isotopes are plotted with different colours: 76Ge, 136Xe, 130Te, 100Mo in green, blue, red and purple, respectively. SNO+ (in red) is shown with increasing amounts of Te (0.5%, 1.5% and 3%). The x-axis is plotted as the expected half-life sensitivity weighted by precisely calculable phase space factors. Different symbols correspond to different nuclear matrix element model calculations.
@@
The feasibility thus exists to scale up a Te-loaded liquid scintillator to eventually reach sensitivity to $m_{\beta\beta}$ in the non-degenerate normal mass ordering range of parameter space. It is for this reason that JUNO and THEIA (Section 6.9) are also considering a large-scale deployment of Te in their liquid scintillator detectors. SNO+ represents a critical first step in this endeavour. Demonstrating how well this approach works and the control of backgrounds during Te purification and loading are objectives, as well as serving as a competitive neutrinoless double beta decay experiment searching with excellent sensitivity in new parameter space, with a window of opportunity over the next 10 years.
@@ 6.4 Exploring sub-GeV Dark Matter in Liquid Argon at SNOLAB: DarkSide20k and Darkside-LowMass
Two purification scenarios are considered for the DSLM target, which will consist of UAr extracted from the URANIA plant in Southwest Colorado. Inspired by DarkSide-50, where a 1400-fold reduction in 39Ar activity was observed in argon extracted from an existing mine and purified at FNAL at 140 g/day, URANIA was built to meet tonne-scale experimental needs and is expected to begin commissioning by summer 2025 with an extraction rate of approximately 250 kg/day. The UAr will then be processed at ARIA using the 350-meter-tall Seruci-1 cryogenic isotopic distillation column to reduce 39Ar down to $\mathcal{O}(10)$ $\mu$Bq/kg, compared to $\mathcal{O}(1)$ mBq/kg achieved in DarkSide-50 UArpath . However, achieving this level of purification requires a 10 kg/day throughput at ARIA, which is feasible for DSLM but unrealistic for DS20k, where ARIA will be used only for standard chemical distillation. Including cosmogenic activation of 39Ar during surface storage and transport to and from North America, DSLM’s final 39Ar content is expected to be approximately 73 $\mu$Bq/kg after one pass through ARIA and 7.3 $\mu$Bq/kg after two DSLMwhitepaper .
@@ 6.7 NEXT: Neutrino Experiment with a Xenon TPC
The program is phased, with the current experiment NEXT-100 next100detector ; next100CDR , currently under operation. Initial data from NEXT-100 has achieved electron drift lifetimes of the order of tens of ms, energy resolution of 5% at 40 keV with ${}^{83\textnormal{m}}$r calibration, and reconstruction of 30 keV x-rays. NEXT-100 is currently taking calibration with a high-energy 232Th source (2.6 MeV gammas).
@@ 6.9 Theia: A Hybrid Neutrino Detector
Long-baseline oscillations $>5\sigma$ for 30% of $\delta_{CP}$ values 524 kt-MW-yr
Galactic Supernova (10 kpc) $<2^{\circ}$ pointing accuracy 25-kt detector
@@
CNO neutrino flux $<$ 10% 62.5 kton-yr
@@ 7.1.5 Positive Ion Enabled (PIE) detector
The diffusion limitation of the previous section suggests that another approach might be useful. Some years ago, Martell et al. Martoff:2000wi pointed out that in most media, but especially in noble gases and liquids, the electrons are very non-thermal and thus have very high diffusion levels. Molecules, on the other hand will remain quite close to thermal. Diffusion is small, even with long drift time because, from the Einstein relations D=$\mu$kT/q, D is small when the mobility is small.
@@ 8 Conclusions: Summary of Requested Capabilities
  • Cryogenic distillation column of sufficient scale and throughput to be of use in both argon and xenon programs which may choose to site themselves at SNOLAB (e.g.  6.5). Installation in Shaft 8 is suggested.
@@
  • Upgrades to the existing chemistry facility, including a dedicated setup for radon removal and noble liquid target purification (e.g. Section  6.5).
  • Large-scale cryogenic storage and recovery systems, for argon and/or xenon (see Sections 6.5, 6.6, 6.7).
@@
  • Preparation for increased demand on the CUTE facility, in the case that more and more new technologies desire a place to test their designs in a well-shielded and cryogenic environment (e.g. Sections 4.457.3). These tests will very likely also lead to scientific output, such as experimental constraints on nuclear matrix element predictions. In addition, the activities at the CUTE facility involving qubits, if successful, may lead to the need to integrate qubits alongside new detector targets with the qubits serving as the active detector elements (see Section 4.6). In general, there is a sense of increasing pressure on the underground CUTE facility from anticipated projects, meriting the consideration of an additional CUTE facilty underground. This could be placed in an existing water tank in the current ladder labs or be part of a new ladder labs if the laboratory were expanded.

### v3

@@ 3.2 The China Jinping Underground Laboratory
[…]
@@ 4.2 HeLIOS – Detecting Gravitational Waves and Dark Matter with Superfluid Helium
Co-Authors: M. Hirschel, J. P. Davis
@@ 4.3 Liquid-Noble Bubble Chambers
[…]
Looking forward to the future, a 1-tonne-yr exposure would allow for a DM search down to the boundary of the argon neutrino fog from 8B solar neutrinos, shown in the right of Fig. 6. A 1 tonne superheated liquid argon volume at 130 K would need dimensions of approximately 1-m-diameter by 1.4-m-high. This volume needs to be pressurized to $\sim 350$ psi after a bubble nucleation, and therefore a pressure vessel capable of housing this fluid would require $\lesssim 0.01$ mBq/kg of primordial isotopes to generate <1<1 background event/yr from fast neutrons. While these are small activities, they may not be so far off from some upper limits, emphasizing the need for precision assay methods in the future.
Critically, synthetic quartz vessels, which are traditionally used to contain the superheated liquid, cannot be produced with active volumes greater than 250 L. This provides an opportunity to investigate other detector materials, such as electropolished metals and acrylic, both of which have been demonstrated in small-scale test chambers Broerman:2022jtj . To limit deadtime from repressurizing the detector after a bubble event ($\sim$30 s/event), surface rates from radiogenic and roughness features would need to be kept <<5 nBq/cm2 to result in <1<1 event/hour. For external shielding, a 9-m-diameter water tank would be sufficient to reduce the fast neutron and Thomson scattering backgrounds from the cavern walls to <1<1 event/year each.
@@ 4.4 Cryogenic Solid-State Detectors: SuperCDMS and Beyond
[…]
@@
    There is the possibility of sharing infrastructure with other future cryogenic solid-state detector projects, e.g. Globolo (see Section 4.7), and synergy with TESSERACT (see Section 3.3.3).
@@ 4.6 Novel Small-Scale Experiments for Light Dark Matter
Co-Authors: Y. Kahn, B. Roach, N. Hoch, A. Williams, E. Marrufo Villalpando, D. Baxter, A. Drlica-Wagner, P. Abbamonte, D. Balut, J. Oh, L. Thompson, D. Johnson, D. Freedman, L. Winslow, C. Blanco, A. Radick, B. Lillard, C. Gao, J. Schütte-Engel, W. Halperin, M. Nguyen, J.W. Scott, J. Foster
@@
4.7 Bolometric Detectors: Precision Beta Decay and Future Neutrinoless Double Beta Decay Searches
@@ 5.2 SNO+ Tellurium Double Beta Decay
SNO+ will be moving forward with the $0\nu\beta\beta$ phase of the experiment. Tellurium will be added to the detector, starting in 2026, in the form of tellurium-butanediol complexes dissolved in the SNO+ liquid scintillator. The SNO+ $0\nu\beta\beta$ search starts after the current pure scintillator phase has completed its main physics program, and after test batches have fully commissioned the two underground tellurium process plants (purification and synthesis). In the first instance, 0.5% Te, by weight, will be the target loading. This amounts to 1.33 tonnes of the isotope 130Te in the detector, a large quantity due to its high natural abundance of 34%. A proposal has been submitted to the CFI 2025 Innovation Fund competition. If it is successful, it would enable SNO+ to triple the amount of Te loaded in the detector to 1.5% Te, reaching 4 tonnes of 130Te deployed for the $0\nu\beta\beta$ search. Figure 8 illustrates the sensitivity that SNO+ 0.5% and SNO+ 1.5% can reach in comparison to other existing, near-future and planned $0\nu\beta\beta$ experiments.
Figure 8: Comparison of $0\nu\beta\beta$ sensitivities for current (Section 3.3), near-term and future proposed (Sections 3.3 and 5.6) experiments. Sensitivity is expressed as $m_{\beta\beta}$. Different isotopes are plotted with different colours: 76Ge, 136Xe, 130Te, 100Mo in green, blue, red and purple, respectively. SNO+ (in red) is shown with increasing amounts of Te (0.5%, 1.5% and 3%). The x-axis is plotted as the expected half-life sensitivity weighted by precisely calculable phase space factors. Different symbols correspond to different nuclear matrix element model calculations.
@@
The feasibility thus exists to scale up a Te-loaded liquid scintillator to eventually reach sensitivity to $m_{\beta\beta}$ in the non-degenerate normal mass ordering range of parameter space. It is for this reason that JUNO and THEIA (Section 5.9) are also considering a large-scale deployment of Te in their liquid scintillator detectors. SNO+ represents a critical first step in this endeavour. Demonstrating how well this approach works and the control of backgrounds during Te purification and loading are objectives, as well as serving as a competitive $0\nu\beta\beta$ experiment searching with excellent sensitivity in new parameter space, with a window of opportunity over the next 10 years.
@@ 5.4 Exploring sub-GeV Dark Matter in Liquid Argon at SNOLAB: DarkSide20k and Darkside-LowMass
[…]
@@ 5.7 NEXT: Neutrino Experiment with a Xenon TPC
The program is phased, with the current experiment NEXT-100 next100detector ; next100CDR , currently under operation. Initial data from NEXT-100 has achieved electron drift lifetimes of the order of tens of ms, energy resolution of 5% at 40 keV with $^{83\textnormal{m}}$Kr calibration, and reconstruction of 30 keV x-rays. NEXT-100 is currently taking calibration with a high-energy 232Th source (2.6 MeV gammas).
@@ 5.9 Theia: A Hybrid Neutrino Detector
Long-baseline oscillations $5\sigma$ for 30% of $\delta_{CP}$ values 524 kt-MW-yr
Galactic Supernova (10 kpc) $2^{\circ}$ pointing accuracy 25-kt detector
@@
CNO neutrino flux << 10% 62.5 kton-yr
@@ 6.1.5 Positive Ion Enabled (PIE) detector
[…]
@@ 7 Conclusions: Summary of Requested Capabilities
  • Cryogenic distillation column of sufficient scale and throughput to be of use in both argon and xenon programs which may choose to site themselves at SNOLAB (e.g.  5.5). Installation in Shaft 8 is suggested.
@@
  • Upgrades to the existing chemistry facility, including a dedicated setup for radon removal and noble liquid target purification (e.g. Section  5.5).
  • Large-scale cryogenic storage and recovery systems, for argon and/or xenon (see Sections 5.5, 5.6, 5.7).
@@
  • Preparation for increased demand on the CUTE facility, in the case that more and more new technologies desire a place to test their designs in a well-shielded and cryogenic environment (e.g. Sections 4.44.76.3). These tests will very likely also lead to scientific output, such as experimental constraints on nuclear matrix element predictions. In addition, the activities at the CUTE facility involving qubits, if successful, may lead to the need to integrate qubits alongside new detector targets with the qubits serving as the active detector elements (see Section 4.6). In general, there is a sense of increasing pressure on the underground CUTE facility from anticipated projects, meriting the consideration of an additional CUTE facilty underground. This could be placed in an existing water tank in the current ladder labs or be part of a new ladder labs if the laboratory were expanded.